Determine the current through each resistor in the circuit .Applying Kirchhoff's rules:(Hint: each resistor does not have the same current flowing through it. Use Kirchoff's loop and junctions laws to find each of the currents. You will need to write equations for more than one loop.) 12V 3Ω 4Ω 5V 2Ω

1
Expert's answer
2021-11-04T16:20:10-0400

Current in the resistors connected in parallel to the source of 12 V (3 loops: V-R1, V-R2, V-R3)


"I_1=\frac V{R_1}=4\text{ A},\\\space\\
I_2=\frac V{R_2}=2.4\text{ A},\\\space\\
I_3=\frac V{R_2}=6\text{ A}."

The total current (in the cource) by Kirchhoff's current law:


"I=I_1+I_2+I_3=12.4\text{ A}."